OpenStudy (anonymous):

How do you solve? r +8 over r = -9

OpenStudy (anonymous):

what is meant by over r = -9??

OpenStudy (anonymous):

fraction 8 - r

OpenStudy (anonymous):

8/ r and then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do you have to find the common denominator

OpenStudy (anonymous):

ur question still not clear

OpenStudy (anonymous):

r+ 8/r =-9 trying to figure out what r is

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

r + 8 = -9r r +9r = -8 10r = -8 r = -8/10 = -4/5

OpenStudy (anonymous):

thanks